# Spin Alignment of Vector Mesons Induced by Local Spin Density Fluctuations

### v2

@@ III.3 Spin alignment of $K^{*0}$
$K^{*0}$ meson is similar with $\phi$, and the only difference is replacing the composed $s$ quark by $d$ quark, and the spin correlation of $s$ and $\bar{d}$ is of form:
@@
where we have simply set $\langle\delta N_{d}\delta N_{\bar{s}}\rangle=\langle\delta N_{s}\delta N_{\bar{d}}\rangle$. For a non-flavor-mixing-interacting system, the spin polarization of $s$ quark and $d$ quark could be completely independent, and it is obvious that no spin alignment of $K^{*0}$ could arise, even the local spin density for $s$ or $d$ quark is none zero. In other words, the observed spin alignment of $K^{*0}$ requires flavor-mixing axial-vector interaction in this framework:
@@
considering that we only study $\phi$ and $K^{*0}$, $\psi=(d,s)^{T}$, and spin potential $\hat{a}$ is also a matrix in flavor space:
@@
now for the first case:
@@
and the thermal dynamics potential receives contribution from FIG.3(a), but the loops are $d$ and $s$ quark, and the only difference is the quark mass. Then we obtain
@@
which is none-zero and could be either positive or negative, depending on the sign of $G_{A}^{mix}$.
Above we consider the how spin polarization of quark induce spin alignment of vector mesons, and for magnetic moment polarization together with the tensor interaction of quarks, the calculation are totally same, and here we don’t repeat.
@@ IV The axial-vector and tensor interactions
The interactions between quarks are complicated at finite temperature and density, especially in the presence with external field like magnetic field or rotation, where the scalar interaction channel is mostly considered because of its close relation with chiral symmetry breaking and restoration [21, 22]. However, all the interaction channels are supposed to exist in general and we have no conceivable reasons to ignore them. Exploring the existence of the axial-vector interaction $G_{A}(\bar{\psi}\gamma^{\mu}\gamma^{5}\psi)^{2}$ and tensor interaction $G_{T}(\bar{\psi}i\gamma^{1}\gamma^{2}\psi)^{2}$, rather than relying solely on assumptions, could contribute to a clearer understanding of the physical picture. In this regard, we will now consider two potential candidates: the One-Gluon-Exchange and instanton-induced interactions.
@@ IV.1 One-Gluon-Exchange
The NJL-type interactions can be extracted from One-Gluon-Exchange process in QCD. Considering the nonzero gluon condensates $\langle A^{2}\rangle\neq 0$ thus an effective large gluon mass $M_{g}$ arises, and the gluon propagator in Feynman gauge has form of:
@@
and the color indices are ignored. However, the gluon propagator could be quite complicated, and here we consider the contribution from one-quark-loop, which lead to
@@
where $\Pi$ is quark loop, and effectively and simply, the quark loop affects the effective gluon mass:
@@
For low energy process $q^{2}\ll\tilde{M}_{g}^{2}$ where low-momentum gluon is interchanged, the four-fermion current-current interaction can be obtained by replacing the gluon propagator $G_{\mu\nu}$ by a constant:
@@
where the coupling constant $G\sim 1/\tilde{M}_{g}^{2}$. Ignore the quark loop contribution for now and $G\sim 1/M_{g}^{2}$ is a constant, then after Fierz transformation in the Dirac space, we know that scalar as well as the axial-vector interaction channel always shows up [21, 24]:
@@
The equation of Fierz transformation Eq.(50) is obtained after summation over four channel $\mu=0,...,3$ and under the assumption that the coupling constants are same for each channel. Now let’s perform Fierz transformation for each channel of vector interaction:
@@
and in the vacuum, all $G^{\mu}$ are equal to each other then the tensor channels are canceled with each other and this is what we found in Eq.(50), while the coupling constants for $\gamma^{0}\times\gamma^{0}$ and $\gamma^{i}\times\gamma^{i}$ could be different at finite temperature and baryon chemical potential in medium, as a result, the tensor interaction will survive.
On the other hand, the axial-vector interaction channel also receives modification from the medium, and a similar analysis lead to the coupling constant of axial-vector interaction $\gamma^{3}\gamma^{5}\times\gamma^{3}\gamma^{5}$:
@@ IV.2 Instanon-induced interactions
The effective interactions induced by OGE does not mix flavor, which will not contribute to the spin alignment of $K^{*)}$. On the other hand, the effective interactions of quarks can also be induced by the instantons[25, 26], and at different temperature, the ensembles of instantons thus the interactions of quarks are also different. What more important is instanton-induced interactions can be flavor-mixing, which could induce non-zero correlation between $s$ quark and $d$ quark thus result in spin alignment of $K^{*0}$.
@@
As temperature increases and approaches chiral phase transition temperature $T\simeq T_{c}$, which is case in heavy ion collision, the (anti-) instantons goes from random phase to a strongly correlated ”molecular” phase ($I\bar{I}$)[29, 26], and in this case, two kind of effective four-fermion axial-vector interaction can always be generated even in 3-flavor case[30] :
@@
with $\lambda^{a}$ the Gell-Mann matrices, then we obtain the $s$ quark and $d$ quark mixing terms
@@ V Numerical estimation
To show whether the fluctuation of spin density could induce spin alignment that comparable to the experiment measurement and help to explain the data, in this section we will perform numerical calculation and estimate its magnitude. The temperature and baryon chemical potential as functions of collision energy can be extracted in heavy ion collisions [31]: $\mu_{B}=1.477/(1+0.343\sqrt{s})$ and $T=0.158-0.14\mu_{B}^{2}-0.04\mu_{B}^{4}$. And it is found that in collision energy region $7$ 200 GeV, the temperature varies from 0.134 GeV to 0.158GeV, which can be regarded as constant, while baryon chemical potential varies from near zero to $0.4$GeV.
As we have dissuasion about the behavior of coupling constants of tensor and axial-vector interactions from OGE, here we simply make the assumptions:
@@
which can be regarded as a Taylor expansion of $G_{T}/G_{A}$ over small $\mu_{B}$, and $c$ is a free parameter. Besides, none-zero baryon chemical potential would have effect on the spin alignment, however here we ignore it as we didn’t consider it in above argument. On the other hand, the effective quark mass are evaluated in a 3-flavor NJL model[32], and it is found that the effective mass of $s$ is $m_{s}\approx 0.425\text{GeV}$, and $m_{d}\approx 0.184\text{GeV}$ for $d$ quark. Notice that the effect from both interactions are small, here the spin alignment of $\phi$ meson is considered as:
@@ VI Conclusion and Discussion
On the other hand, to find the SP background from mean field is straightforward. However, consider the presence of the axial-vector interaction of quarks, we found that the $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle\neq 0$ for $G_{AV}\gtrsim G_{S}$, which means that global and homogeneous condensate of SP background is not obtained in the mean field approximation. On the other hand, the spin alignment of vector meson induced by SP(MMP) background does not depend on the sign of the background, thus, through the fluctuation of spin density, locally nonzero SP(MMP) background can be obtained, thus $\rho_{00}$ of the vector mesons emitted from local region has deviation from $1/3$. We then calculated $\langle(\bar{\psi}\gamma^{3}\gamma^{5}\psi)^{2}\rangle$ instead of $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$, then nonzero spin density is obtained thus spin alignment of vector mesons.
We also argued the source of interactions of quarks. Generally, all the forms of quark interaction could exist in vacuum and finite temperature, however, we still need to specify how the interaction arises. We find two possible sources, and one is from the One-Gluon-Exchange process which induces axial-vector interaction, and the tensor interaction can also be generated, and both coupling constants are supposed to be a function of temperature as well as the other conditions. Another source of interaction is induced by instanton, which could gives the tensor interaction at low temperature and axial-vector interaction around and above $T_{c}$. The instanton-induced interactions are also flavor mixed, which is absent in OGE, and could explain the spin alignment of $K^{*0}$.
This mechanism could contribute to the spin alignment at central collision, but also to the cases of non-central collision. On the other hand, the coupling constants of axial-vector and tensor interaction induced by OGE or instantons are functions of the angular velocity, magnetic field as well as the baryon chemical potential, which depends on the centrality and collision energy.

### v3

@@ III.3 Spin alignment of $K^{*0}$
$K^{*0}$ meson is similar to $\phi$, and the only difference is replacing the composed $s$ quark by $d$ quark, and the spin correlation of $s$ and $\bar{d}$ is of form:
@@
where we have simply set $\langle\delta N_{d}\delta N_{\bar{s}}\rangle=\langle\delta N_{s}\delta N_{\bar{d}}\rangle$. For a non-flavor-mixing-interacting system, the spin polarization of $s$ quark and $d$ quark could be completely independent, and it is obvious that no spin alignment of $K^{*0}$ could arise, even the local spin density for $s$ or $d$ quark is non-zero. In other words, the observed spin alignment of $K^{*0}$ requires flavor-mixing axial-vector interaction in this framework:
@@
considering that we only study $\phi$ and $K^{*0}$, thus, $\psi=(d,s)^{T}$, and spin potential $\hat{a}$ is also a matrix in flavor space:
@@
then the spin correlation between $d$ and $\bar{s}$ is:
@@
and the thermodynamic potential receives contribution from FIG.3(a), but the loops are $d$ and $s$ quark, and the only difference is the quark mass. Then we obtain
@@
which is non-zero and could be either positive or negative, depending on the sign of $G_{A}^{mix}$.
Above we consider how spin polarization of quarks induce spin alignment of vector mesons, and for magnetic moment polarization together with the tensor interaction of quarks, the calculation is totally same, and here we don’t repeat.
@@ IV The axial-vector and tensor interactions
The interactions between quarks are complicated at finite temperature and density, especially in the presence with external fields like magnetic field or rotation, where the scalar interaction channel is mostly considered because of its close relation with chiral symmetry breaking and restoration [21, 22]. However, all the interaction channels are supposed to exist in general and we have no conceivable reasons to ignore them. Exploring the existence of the axial-vector interaction $G_{A}(\bar{\psi}\gamma^{\mu}\gamma^{5}\psi)^{2}$ and tensor interaction $G_{T}(\bar{\psi}i\gamma^{1}\gamma^{2}\psi)^{2}$, rather than relying solely on assumptions, could contribute to a clearer understanding of the physical picture. In this regard, we will now consider two potential candidates: the One-Gluon-Exchange and instanton-induced interactions.
@@ IV.1 One-Gluon-Exchange
The NJL-type interactions can be extracted from One-Gluon-Exchange process in QCD. Considering the non-zero gluon condensates $\langle A^{2}\rangle\neq 0$ thus an effective large gluon mass $M_{g}$ arises, and the gluon propagator in Feynman gauge has form of:
@@
and the color indices are ignored. However, the gluon propagator could be quite complicated, and here we consider the contribution from one-quark-loop, which leads to
@@
and $\Pi$ is quark loop polarization function, which simply affects the effective gluon mass:
@@
For low energy process $q^{2}\ll\tilde{M}_{g}^{2}$ where low-momentum gluon is interchanged between quarks, the four-fermion current-current interaction can be obtained by replacing the gluon propagator $G_{\mu\nu}$ by a constant:
@@
where the coupling constant $G\sim 1/\tilde{M}_{g}^{2}$. Ignore the quark loop contribution for now and $G\sim 1/M_{g}^{2}$ is a constant, then after Fierz transformation in the Dirac space, we know that scalar as well as the axial-vector interaction channels always show up [21, 26]:
@@
The equation of Fierz transformation Eq.(50) is obtained after summation over four channels $\mu=0,...,3$ and under the assumption that the coupling constants are same for each channel. Now let’s perform Fierz transformation for each channel of vector interaction:
@@
and in the vacuum, all $G^{\mu}$ are equal to each other then the tensor channels are canceled with each other and this is what we found in Eq.(50), however, the coupling constants for $\gamma^{0}\times\gamma^{0}$ and $\gamma^{i}\times\gamma^{i}$ could be different at finite temperature and baryon chemical potential in medium, as a result, the tensor interaction will survive.
On the other hand, the axial-vector interaction channel also receives modification from the medium, and a similar analysis leads to the coupling constant of axial-vector interaction $\gamma^{3}\gamma^{5}\times\gamma^{3}\gamma^{5}$:
@@ IV.2 Instanon-induced interactions
The effective interactions induced by OGE does not mix flavors, which will not contribute to the spin alignment of $K^{*}$. On the other hand, the effective interactions of quarks can also be induced by the instantons[27, 28], and at different temperature, the ensembles of instantons thus the interactions of quarks are also different. What more important is instanton-induced interactions can be flavor-mixing, which could induce non-zero correlation between $s$ quark and $d$ quark thus result in spin alignment of $K^{*0}$.
@@
As temperature increases and approaches chiral phase transition temperature $T\simeq T_{c}$, which happens in heavy ion collision, the (anti-) instantons goes from random phase to a strongly correlated ”molecular” phase ($I\bar{I}$)[31, 28], and in this case, two kind of effective four-fermion axial-vector interactions can always be generated even in 3-flavor case[32] :
@@
with $\lambda^{a}$ the Gell-Mann matrices, then we obtain the $s$ quark and $d$ quark mixing term from the second one:
@@ V Numerical estimation
To show whether the fluctuation of spin density could induce spin alignment that comparable to the experiment measurement and help to explain the data, in this section we will perform numerical calculation and estimate its magnitude. The temperature and baryon chemical potential as functions of collision energy can be extracted from heavy ion collisions [33]: $\mu_{B}=1.477/(1+0.343\sqrt{s})$ and $T=0.158-0.14\mu_{B}^{2}-0.04\mu_{B}^{4}$. And it is found that in collision energy region $7$ 200 GeV, the temperature varies from 0.134 GeV to 0.158GeV, which can be regarded as a constant, while baryon chemical potential varies from near zero to $0.4$GeV.
As we have discussed about the behavior of coupling constants of tensor and axial-vector interactions from OGE, here we simply make the assumptions:
@@
which can be regarded as a Taylor expansion of $G_{T}/G_{A}$ over small $\mu_{B}$, and $c$ is a free parameter. Besides, non-zero baryon chemical potential would have effect on the spin alignment, however here we ignore it as we didn’t consider it in above argument. On the other hand, the effective quark mass are evaluated in a 3-flavor NJL model[34], and it is found that the effective mass of $s$ is $m_{s}\approx 0.425\text{GeV}$, and $m_{d}\approx 0.184\text{GeV}$ for $d$ quark. Notice that the effect from both interactions are small, here the spin alignment of $\phi$ meson is considered as:
@@ VI Conclusion and Discussion
On the other hand, calculating the SP background from mean field is straightforward. However, consider the presence of the axial-vector interaction of quarks, we found that the $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle\neq 0$ for $G_{A}\gtrsim G_{S}$, which means that global and homogeneous condensate of SP background can not obtained in the mean field approximation. On the other hand, the spin alignment of vector meson induced by SP(MMP) background does not depend on the sign of the background, thus, through the fluctuation of spin density, locally non-zero SP(MMP) background can be obtained, thus $\rho_{00}$ of the vector mesons emitted from local region has deviation from $1/3$. We then calculated $\langle(\bar{\psi}\gamma^{3}\gamma^{5}\psi)^{2}\rangle$ instead of $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$, then non-zero spin density is obtained thus spin alignment of vector mesons appears.
We also argued the source of interactions of quarks. Generally, all the forms of quark interaction could exist in vacuum and finite temperature, however, we still need to specify how the interaction arises. We find two possible sources, and one is from the One-Gluon-Exchange process which induces both axial-vector and tensor interactions, and both coupling constants are supposed to be functions of temperature as well as the other conditions. Another source of interaction is instanton, which could gives the tensor interaction at low temperature and axial-vector interaction around and above $T_{c}$. The instanton-induced interactions are also flavor-mixed, which is absent in OGE, and could explain the spin alignment of $K^{*0}$.
This mechanism could contribute to the spin alignment at central collision, as well as to the cases of non-central collision. On the other hand, the coupling constants of axial-vector and tensor interaction induced by OGE or instantons are functions of the angular velocity, magnetic field as well as the baryon chemical potential, which depends on the centrality and collision energy.